OpenStudy (anonymous):

Find three consecutive even integers such that the sum of the smallest and twice the second Is twenty more than the third

OpenStudy (anonymous):

you can get these equation from the question: let x,y,z be the 3 consecutive numbers. x + y -z = 20, z-y = 1 y-z = 1 then try solving them: y = z - 1 x = y - 1, x = (z-1) -1 then put them in the first equation: (z-2) + (z-1) - z = 20 z - 3 = 20 z = 23 so you get the three numbers which are 21,22 and 23